Find the equation of the plane passing through the points (1,1,2) and (2,2,2) and perpendicular to the plane 6x2y+2z = 9.

Grade:Upto college level

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Please find the answer to your question below
the equation of the plane passing through the points (1,1,2) and (2,2,2) and perpendicular to the plane 6x2y+2z = 9 is
\begin{vmatrix} x-1 &y-1 &z-2 \\ 1& 1 &0 \\ 6& -2 &2 \end{vmatrix}=0
which gives x-y-4z+8=0